Grade 11General Physics

How it`s p= mat???? In dimensional formula Q-1 If linear momentum p , acceleration a, and time t , are taken as fundamental quantities then dimensional formula for mass is ??

Joy Mattu
8 Years agoGrade 11

1 Answer

Eshan
8 Years ago
If accelerationaof a body is constant, the force acting on the bodyF=\dfrac{\Delta p}{\Delta t}=ma

\implies \Delta p=ma\Delta t

Hence\implies p=mat

From the formula,m=\dfrac{p}{at}
Hence dimentional formula of mass=[M]=\dfrac{[p]}{[A][T]}